OpenStudy (anonymous):

Is the horizontal asymptote of h(x)=-(1/4x), y=0?

OpenStudy (anonymous):

@Nanalew @sourwing @zzr0ck3r

OpenStudy (anonymous):

@mathmale

OpenStudy (zzr0ck3r):

sorry something came up

OpenStudy (zzr0ck3r):

\(-\frac{1}{4x}\)?

OpenStudy (anonymous):

Completely understandable. I actually got this one, so you think we can do a different one?

OpenStudy (zzr0ck3r):

yes, H-asymptote is y = 0, and V-asymptote is x = 0

OpenStudy (zzr0ck3r):

sure just post it in here

OpenStudy (anonymous):

Well I actually have two. One is quick. Can there be a radical in the numerator of a rational function?

OpenStudy (zzr0ck3r):

you can have a radical anywhere

OpenStudy (anonymous):

Alright, thank you very much.

OpenStudy (zzr0ck3r):

some teachers just prefer you do change the form such that the radical is not in the denominator, but there is actually nothing wrong with it.

OpenStudy (zzr0ck3r):

its sort of like if you answer a question with x = 4/2 , the answer is not wrong but your teacher wants you to write x = 2

OpenStudy (anonymous):

One second this equation stuff takes forever

OpenStudy (anonymous):

\[\frac{ x ^{3}-5x+4}{ 6x-9 }\]

OpenStudy (anonymous):

Is that a rational function?

OpenStudy (anonymous):

With a domain of x=/=(9/6)?

OpenStudy (zzr0ck3r):

yes

OpenStudy (anonymous):

Alright, that's all. Thanks for everything. If I have any other questions I'll be back

OpenStudy (zzr0ck3r):

word gl